OpenStudy (anonymous):

Index law help please! Find x: 36^2x+1 / 216^x-2 = 1/6 In the textbook it says the answer is -9, but I got something completely different. Here is what I did, please tell me what I'm doing wrong, thankyou! (6^2)^2x+1/(6^3)^x-2 = 1/6 6^4x+2 / 6^3x-6 = 1/6 4x+2 / 3x-6 = 1/6 (here i cancelled the 6's-not right I don't think.) 1x-4 = 1/6 (used index law of a^m/ a^n = a^m-n) therefore, x=-10 Please & thankyou

hartnn (hartnn):

you can't cancel the 6's ....as u suspected

OpenStudy (anonymous):

yes don't cancel the six

hartnn (hartnn):

1/6 = 6^{-1} then compare the exponents

OpenStudy (anonymous):

Thankyou so much, so what happens instead of that?

OpenStudy (anonymous):

a^m/ a^n =a^m-n

hartnn (hartnn):

fitst simplify 6^(....)/6^(...) using that ^

OpenStudy (anonymous):

so 6 ^4x-2-(3x-6)

hartnn (hartnn):

\(\huge 6^{(4x+2)-(3x-6)}=...?\)

OpenStudy (anonymous):

therefore 6^(4x-2-3x+6)=1/6

OpenStudy (anonymous):

6^(x+4)=1/6

OpenStudy (anonymous):

Thank-you so much! I get it now :)

OpenStudy (anonymous):

1/6 = 6^-1

hartnn (hartnn):

\(\huge 6^{(4x+2)-(3x-6)}=6^{-1} \\ \huge \implies (4x+2)-(3x-6)=-1 \)

OpenStudy (anonymous):

x+4 =-1 therefore x =3

hartnn (hartnn):

actually x+8 = -1

hartnn (hartnn):

its 4x+2, not 4x-2 :)

OpenStudy (anonymous):

So I got 3 for my answer, but the textbook says -9...

OpenStudy (yttrium):

it's -9 @student.1 since you're gonna distribute the negative sign at 3x-6

hartnn (hartnn):

how u got 3 ?

OpenStudy (yttrium):

so what will happen is that it'll become 4x+2-3x+6 = -1 x+8 = -1 Therefore, x =-9

OpenStudy (anonymous):

OH thanks